OpenStudy (anonymous):

m+3(m-4)=16 Steps i need also on how to get this.

OpenStudy (karatechopper):

wel first distribute the 4 and get rid of the parenthesis

OpenStudy (karatechopper):

my mistake i mean distribute the 3

OpenStudy (anonymous):

m + 3m - 12 = 16 4m - 12 = 16 4m = 16 + 12 = 28 m = 28/4 = 7

OpenStudy (karatechopper):

m+3m-12=16 llike terms combine 4m-12=16 add 12 to both sides 4m=28 divide buy 4 on both sides 4m/4=28/4 m=7

OpenStudy (anonymous):